# Holographic Tensor Networks as Tessellations of Geometry

### v1

@@ Appendix A The RT formula in the factorized PEE tensor network
Before making a comparison between the factorized PEE tensor network with the HaPPY code [46], we first introduce some preliminary about isometric tensor for the section to be self-contained.
@@
Figure 4: (a) The three type of PEE threads, denoted by $\mathcal{C}^{A}$, $\mathcal{C}^{\bar{A}}$ and $\mathcal{C}^{A\bar{A}}$, are labeled by gray, orange and blue solid simi-circles respectively, while the RT surface $\gamma_{A}$ is drawn as the red dashed line; (b) The circuit interpretation of the factorized PEE tensor network. Each PEE thread individually gives a unitary map, and the RT surface $\gamma_{A}$ only cuts the threads in $\mathcal{C}^{A\bar{A}}$ once. The graph can be compared with Fig.19 of [46] where the “circuit” is built from the net effect of perfect tensors of certain bulk subregions as unitary gates.
It is also useful to derive the RT formula using the arguments in [46] for more generic tensor network models. To be specific, let us cut the tensor network open along a homologous surface $\mathcal{C}$ whose boundary matches the boundary of $A$, such that the tensor network is divided into two parts $\mathcal{P}$ and $\mathcal{Q}$, which are maps from $\mathcal{C}$ to $A$ and from $\mathcal{C}$ to $\bar{A}$ respectively. The boundary state can be written as
@@ Appendix B HaPPY-like PEE tensor network
As was shown in [46], a bulk tensor network built from perfect tensors also satisfy the entropy bound (32), and the inequality is saturated if and only if the two tensors $\mathcal{P}$ and $\mathcal{Q}$ obtained by cutting the whole bulk tensor along $\gamma_{A}$ are isometric. Since the graph of our network is the PEE network, again we have $\mathcal{N}(\gamma_{A})={\text{Area}(\gamma_{A})}/{4G}$. In order to reproduce the RT formula, we only need to demonstrate that the mapping $\mathcal{P}$ is an isometry that maps $\gamma_{A}$ to $A$ and $\mathcal{Q}$ is an isometry that map $\gamma_{A}$ to $\bar{A}$.
Now we build the isometric map from $\gamma_{A}$ to $A$ for our HaPPY-like PEE tensor network using an analogous greedy algorithm proposed in [46]. The procedure is decomposed into the following steps:

### v2

@@ Appendix A The RT formula in the factorized PEE tensor network
Before making a comparison between the factorized PEE tensor network with the HaPPY code Pastawski:2015qua , we first introduce some preliminary about isometric tensor for the section to be self-contained.
@@
Figure 4: (a) The three type of PEE threads, denoted by $\mathcal{C}^{A}$, $\mathcal{C}^{\bar{A}}$ and $\mathcal{C}^{A\bar{A}}$, are labeled by gray, orange and blue solid simi-circles respectively, while the RT surface $\gamma_{A}$ is drawn as the red dashed line; (b) The circuit interpretation of the factorized PEE tensor network. Each PEE thread individually gives a unitary map, and the RT surface $\gamma_{A}$ only cuts the threads in $\mathcal{C}^{A\bar{A}}$ once. The graph can be compared with Fig.19 of Pastawski:2015qua where the “circuit” is built from the net effect of perfect tensors of certain bulk subregions as unitary gates.
It is also useful to derive the RT formula using the arguments in Pastawski:2015qua for more generic tensor network models. To be specific, let us cut the tensor network open along a homologous surface $\mathcal{C}$ whose boundary matches the boundary of $A$, such that the tensor network is divided into two parts $\mathcal{P}$ and $\mathcal{Q}$, which are maps from $\mathcal{C}$ to $A$ and from $\mathcal{C}$ to $\bar{A}$ respectively. The boundary state can be written as
@@ Appendix B HaPPY-like PEE tensor network
As was shown in Pastawski:2015qua , a bulk tensor network built from perfect tensors also satisfy the entropy bound (32), and the inequality is saturated if and only if the two tensors $\mathcal{P}$ and $\mathcal{Q}$ obtained by cutting the whole bulk tensor along $\gamma_{A}$ are isometric. Since the graph of our network is the PEE network, again we have $\mathcal{N}(\gamma_{A})={\text{Area}(\gamma_{A})}/{4G}$. In order to reproduce the RT formula, we only need to demonstrate that the mapping $\mathcal{P}$ is an isometry that maps $\gamma_{A}$ to $A$ and $\mathcal{Q}$ is an isometry that map $\gamma_{A}$ to $\bar{A}$.
Now we build the isometric map from $\gamma_{A}$ to $A$ for our HaPPY-like PEE tensor network using an analogous greedy algorithm proposed in Pastawski:2015qua . The procedure is decomposed into the following steps: